In the country of Economica, the total labor force consists of 5,000 workers. 200 of these workers are unemployed, and thus the current unemployment rate is 4.0%. This year, 100 workers will lose their jobs, 85 workers will return to the labor force after an absence, 50 workers will voluntarily leave their jobs, and 250 new workers will enter the labor force. At the same time, 415 workers will find jobs, and 150 workers will retire or leave the labor force for other reasons. Thus, at the end of he year he labor force wil consist o workers, and the unemployment rate be % R ound your responses to one dean a placeJims favorite breakfast is two eggs, 1/5 lb. of bacon, 1/7 lb. of white toast, 1/6 can (16 ounce can) of frozen concentrate orange juice, and 1 lb. of grapefruit. The following table contains 1991 and 2001 prices for these five items. Using 1991 as the base year, compute the price index for Jims breakfast for 2001 Item 1991 Price 2001 Price One dozen eggs One pound bacon One pound bread One can orange juice 1.98 One pound grapefruit $1.08 $2.15 $0.63 $1.02 $2.75 0.98 $1.86 S0.55 $0.57 OA. 89.53 ОС. 93.02 O D. 107.5

Help me with these two questions !! I would love an explanation as well
Answer #1

Labour force consists of all employed people and also the unemployed people who are actively looking for work within age 16 to 64.

Initially there are 5000 people in labor force, out of which 200 were unemployed. So total employed becomes 4800.

This year 100 lose their jobs and 50 will voluntarily leave the job and 85 return back to the labour force and 250 new people will join the labor force and 150 workers will  leave the labor force and 415 will get jobs.

So the new labor force will be = 5000+85+250-150 = 5185

And number of unemployed people now is = 200+100+50+85+250-415 = 270

The new Unemployment rate = 270/5185 = 0.052 = 5.2%

2)Price of 2 eggs in 1991 = (1.08/12)*2 = $0.18

Price of 2 eggs in 2001= (1.02/12)*2 = $0.17

Price of 1/5 pounds of bacon in 1991= 2.15/5 = $0.43

Price of 1/5 pounds of bacon in 2001= 2.75/5 = $0.55

Price of 1/7 pounds of white toast in 1991= 0.63/7 = $0.09

Price of 1/7 pounds of white toast in 2001= 0.98/7 = $0.14

Price of 1/6 ounce of orange juice in 1991=1.98/6 = $0.33

Price of 1/6 ounce of orange juice in 2001 = 1.86/6 = $0.31

Price of 1 pound of grapefruit in 1991 = $0.57

Price of 1 pound of grapefruit in 2001 = $0.55

So price index for Jim’s breakfast in 2001 with 1991 as base year=

(Prices in 2001/Prices in 1991)*100 = ((0.18+0.43+0.09+0.33+0.57)/(0.17+0.55+0.14+0.31+0.55)*100)

= (1.6/1.72)*100 = 93.02.

So option C is the correct answer.